# Unlearning Isn’t Deletion: Investigating Reversibility of Machine Unlearning in LLMs

### v1

@@ A.6 Detailed Analysis Results
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
@@
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
@@
(a) $N=6$, Layer 28
(b) $N=50$, Layer 28
(c) $N=100$, Layer 28
(d) $N=6$, Layer 22
(e) $N=50$, Layer 22
(f) $N=100$, Layer 22
(g) $N=6$, Layer 13
(h) $N=50$, Layer 13
(i) $N=100$, Layer 13
(j) $N=6$, Layer 4
(k) $N=50$, Layer 4
(l) $N=100$, Layer 4
(m) $N=6$, Layer 1
(n) $N=50$, Layer 1
(o) $N=100$, Layer 1
@@
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
@@
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
@@
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
@@
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
@@
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
@@
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 24
(e) LR=$5\times 10^{-6}$, Layer 24
(f) LR=$3\times 10^{-5}$, Layer 24
(g) LR=$3\times 10^{-6}$, Layer 12
(h) LR=$5\times 10^{-6}$, Layer 12
(i) LR=$3\times 10^{-5}$, Layer 12
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
@@
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 24
(e) LR=$5\times 10^{-6}$, Layer 24
(f) LR=$3\times 10^{-5}$, Layer 24
(g) LR=$3\times 10^{-6}$, Layer 12
(h) LR=$5\times 10^{-6}$, Layer 12
(i) LR=$3\times 10^{-5}$, Layer 12
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
@@
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 24
(e) LR=$5\times 10^{-6}$, Layer 24
(f) LR=$3\times 10^{-5}$, Layer 24
(g) LR=$3\times 10^{-6}$, Layer 12
(h) LR=$5\times 10^{-6}$, Layer 12
(i) LR=$3\times 10^{-5}$, Layer 12
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
@@
(d) Simple (input data = forget set Layer), 4
(e) Simple (input data = retain set Layer), 4
@@
A.6.1 Principal Component Analysis: Similarity and Shift
Across the same hyper-parameter grid, Figures 7– 13 plot the PCA–Shift trajectories—layer-wise displacements of activation centroids along the first two principal directions. For GA-based objectives the pattern mirrors their Similarity curves: as LR rises the orange triangles (unlearned) shoot far from the blue circles (original), especially in deeper layers, and the green squares (relearned) fail to return, producing long grey rays that diagnose irreversible drift. With GA+GD the spread is smaller but still widens sharply at $3{\times}10^{-5}$, confirming that doubling the loss term does not prevent global collapse.
NPO and NPO+KL behave differently. Even at aggressive LR the shifts remain tightly clustered—most layers move $<\!100$ units on pc1 and almost none on pc2—and green squares consistently fall back a little onto the original line segment. RLabel shows an intermediate picture: early layers barely move, while late layers fan out as LR or $N$ grow; nonetheless the rays shorten markedly after relearning, indicating that most distortion is still recoverable.
Task complexity amplifies divergence. On Qwen-2.5-7B the GA rays explode more quickly (Figure 13c,f,i), spanning thousands of pc1 units and driving pc2 to extreme negative values—the hallmark of a high-order, multi-layer perturbation predicted by our theory in Section 4. Conversely, NPO’s clusters expand only modestly, even under the same LR, and contract once relearning is applied.
Taken together, PCA-Shift complements PCA-Similarity: Similarity captures angular mis-alignment, while Shift quantifies translational drift. Their joint reading confirms that GA (with or without GD/KL) is prone to large, irreversible representational displacements, whereas NPO variants and, to a lesser extent, RLabel confine shifts to a regime that remains correctable—consistent with the reversible versus irreversible forgetting boundary observed in our utility experiments.
A.6.2 Centered Kernel Alignment Analysis
Figures 15– 17 report layer-wise linear CKA between the original model and its unlearned / relearned counterparts. Across both Yi-6B and Qwen-2.5-7B, GA again stands out: as LR or $N$ grows its CKA curve close to zero in the final third of the network and never returns, revealing a deep sub-space fracture that matches the irreversible PCA trends. GA+GD and GA+KL modestly attenuate this dip, but still fail to restore full alignment after relearning.
By contrast, NPO and NPO+KL keep CKA higher GA through almost all layers—even under $3{\times}10^{-5}$ or $N{=}100$—and relearning lifts the few layers back to baseline, confirming their perturbations are lower than GA series. RLabel occupies an intermediate position: as LR or N increases, its CKA curve drops rapidly—mirroring GA’s behavior—and ultimately exhibits irreversible forgetting.
Task complexity does not change the ordering but amplifies the gaps: on the math-heavy Qwen benchmark GA’s tail layers fall to almost zero at high LR, whereas NPO keeps higher than GA. Taken together with the PCA-Shift results, CKA shows that only GA-style objectives consistently destroy the encoder–decoder sub-space, while NPO families maintain higher stability than GA series and RLabel induces a moderate, recoverable tilt.
A.6.3 Fisher Information Analysis
[…]

### v2

@@ A.5.3 Fisher Information Analysis
[…]
Figure 19: FIM for GA Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
[…]
Figure 20: FIM for GA+GD Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
[…]
Figure 21: FIM for GA+KL Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
[…]
Figure 22: FIM for NPO Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
[…]
Figure 23: FIM for NPO+KL Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
[…]
Figure 24: FIM for Rlable Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
(a) Simple $N=6$, Layer 28
(b) Simple $N=50$, Layer 28
(c) Simple $N=100$, Layer 28
(d) Simple $N=6$, Layer 22
(e) Simple $N=50$, Layer 22
(f) Simple $N=100$, Layer 22
(g) Simple $N=6$, Layer 13
(h) Simple $N=50$, Layer 13
(i) Simple $N=100$, Layer 13
(j) Simple $N=6$, Layer 4
(k) Simple $N=50$, Layer 4
(l) Simple $N=100$, Layer 4
(m) Simple $N=6$, Layer 1
(n) Simple $N=50$, Layer 1
(o) Simple $N=100$, Layer 1
@@
(a) Simple $N=6$, Layer 31
(b) Simple $N=50$, Layer 31
(c) Simple $N=100$, Layer 31
(d) Simple $N=6$, Layer 28
(e) Simple $N=50$, Layer 28
(f) Simple $N=100$, Layer 28
(g) Simple $N=6$, Layer 22
(h) Simple $N=50$, Layer 22
(i) Simple $N=100$, Layer 22
(j) Simple $N=6$, Layer 13
(k) Simple $N=50$, Layer 13
(l) Simple $N=100$, Layer 13
(m) Simple $N=6$, Layer 4
(n) Simple $N=50$, Layer 4
(o) Simple $N=100$, Layer 4
(p) Simple $N=6$, Layer 1
(q) Simple $N=50$, Layer 1
(r) Simple $N=100$, Layer 1
@@
(a) Simple $N=6$, Layer 31
(b) Simple $N=50$, Layer 31
(c) Simple $N=100$, Layer 31
(d) Simple $N=6$, Layer 28
(e) Simple $N=50$, Layer 28
(f) Simple $N=100$, Layer 28
(g) Simple $N=6$, Layer 22
(h) Simple $N=50$, Layer 22
(i) Simple $N=100$, Layer 22
(j) SimpleN=6N=6, Layer 13
(k) Simple $N=50$, Layer 13
(l) Simple $N=100$, Layer 13
(m) Simple $N=6$, Layer 4
(n) Simple $N=50$, Layer 4
(o) Simple $N=100$, Layer 4
(p) Simple $N=6$, Layer 1
(q) Simple $N=50$, Layer 1
(r) Simple $N=100$, Layer 1
@@
(a) Simple $N=6$, Layer 31
(b) Simple $N=50$, Layer 31
(c) Simple $N=100$, Layer 31
(d) SimpleN=6N=6, Layer 28
(e) Simple $N=50$, Layer 28
(f) Simple $N=100$, Layer 28
(g) Simple $N=6$, Layer 22
(h) Simple $N=50$, Layer 22
(i) Simple $N=100$, Layer 22
(j) Simple $N=6$, Layer 13
(k) Simple $N=50$, Layer 13
(l) Simple $N=100$, Layer 13
(m) Simple $N=6$, Layer 4
(n) Simple $N=50$, Layer 4
(o) Simple $N=100$, Layer 4
(p) Simple $N=6$, Layer 1
(q) Simple $N=50$, Layer 1
(r) Simple $N=100$, Layer 1
@@
(a) Simple $N=6$, Layer 31
(b) Simple $N=50$, Layer 31
(c) Simple $N=100$, Layer 31
(d) Simple $N=6$, Layer 28
(e) Simple $N=50$, Layer 28
(f) Simple $N=100$, Layer 28
(g) Simple $N=6$, Layer 22
(h) Simple $N=50$, Layer 22
(i) Simple $N=100$, Layer 22
(j) Simple $N=6$, Layer 13
(k) Simple $N=50$, Layer 13
(l) SimpleN=100N=100, Layer 13
(m) Simple $N=6$, Layer 4
(n) Simple $N=50$, Layer 4
(o) Simple $N=100$, Layer 4
(p) Simple $N=6$, Layer 1
(q) Simple $N=50$, Layer 1
(r) SimpleN=100N=100, Layer 1
@@
(a) Simple $N=6$, Layer 31
(b) Simple $N=50$, Layer 31
(c) Simple $N=100$, Layer 31
(d) Simple $N=6$, Layer 28
(e) Simple $N=50$, Layer 28
(f) Simple $N=100$, Layer 28
(g) Simple $N=6$, Layer 22
(h) Simple $N=50$, Layer 22
(i) Simple $N=100$, Layer 22
(j) Simple $N=6$, Layer 13
(k) Simple $N=50$, Layer 13
(l) Simple $N=100$, Layer 13
(m) Simple $N=6$, Layer 4
(n) Simple $N=50$, Layer 4
(o) Simple $N=100$, Layer 4
(p) Simple $N=6$, Layer 1
(q) Simple $N=50$, Layer 1
(r) Simple $N=100$, Layer 1
@@
[…]
Figure 31: FIM for GA Across Layers. All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.
[…]
Figure 32: FIM for NPO Across Layers. All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.
[…]
Figure 33: FIM for Rlable Across Layers. All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.
@@
(d) Simple (input data = forget set), Layer 4
(e) Simple (input data = retain set), Layer 4